OpenStudy (anonymous):

Use the definitions of the trigonometric ratios for a right triangle to derive a cofunction identity for tan(90° - A). Show your steps.

OpenStudy (anonymous):

\[Tan(90° - A) = \frac{ Sin(90°-A) }{ Cos(90°-A) }=\frac{ Cos(A) }{ Sin(A) } = Cot(A)\]